# SHyPar: A Spectral Coarsening Approach to Hypergraph Partitioning

### v2

@@ II-A7 Hypergraph conductance
A hypergraph $H=(V,E,w)$ consists of a vertex set $V$ and a set of hyperedges $E$ with unit weight $w=1$. The degree of a vertex $d_{v}$ is defined as: $d_{v}:=\Sigma_{e\in E:v\in e}w(e)$, where $w(e)$ represents the weight of each hyperedge. The volume of a node set $S\subseteq V$ in the hypergraph is defined as: $vol(S):=\Sigma_{v\in S}d_{v}$. The conductance of a subset $S$ within the hypergraph is then calculated as:
@@ II-B3 Partitioning objectives
Hypergraph partitioning extends the concept of graph partitioning. Its objective is to distribute the vertex set into multiple disjoint subsets while minimizing a specified cut metric and adhering to certain imbalance constraints. The process of dividing into two subsets is known as bipartitioning, whereas dividing into multiple subsets, typically referred to as $k$-way partitioning, involves partitioning into $k$ parts. More formally, consider a hypergraph $H=(V,E,w)$, where $k$ is a positive integer (with $k\geq 2$) and $\epsilon$ is a positive real number (where $\epsilon\leq\frac{1}{k}$). The objective of $k$-way balanced hypergraph partitioning is to divide $V$ into $k$ disjoint subsets $S=\{V_{0},V_{1},…,V_{k-1}\}$ such that:
@@ III SHyPar: Hypergraph Partitioning via Spectral Coarsening
To address the limitations of existing hypergraph coarsening methods that rely on simple heuristics, we proposed a theoretically sound and practically efficient framework for hypergraph coarsening. Specifically, we proposed a two-phase spectral hypergraph coarsening scheme based on the recent research on spectral hypergraph clustering [46, 47]. Phase A utilizes spectral hypergraph coarsening (HyperEF) to decompose a given hypergraph into smaller node partitions with bounded effective-resistance diameters [47]. This is followed by Phase B, which guides the coarsening stage using a flow-based community detection method (HyperSF) aimed at minimizing the ratio cut [46]. Next, we exploit the proposed two-phase spectral hypergraph coarsening method for multilevel hypergraph partitioning: the prior heuristic hypergraph coarsening schemes will be replaced by the proposed spectral coarsening methods to create a hierarchy of coarser hypergraphs that can preserve the key structural properties of the original hypergraph.
III-A Resistance-Based Hypergraph Clustering (Phase A)
The existing coarsening algorithms, contract vertices at each level of the hierarchy. The primary method involves contracting each vertex with the best neighboring node. This is commonly done by using rating functions to identify and contract highly connected vertices. However, these determinations often rely solely on the weights and sizes of the hyperedges. Rating functions, such as those described in Eq. (9), based on hyperedge size are limited to the local structural properties of the hypergraph and do not account for its global structure. In contrast, the effective-resistance diameter provides a more comprehensive criterion, which considers the global structure. Through an example, we show that hyperedge size-based score functions are not optimal, occasionally disrupting the global structure of the hypergraph. For instance, in scenarios where the hypergraph contains a bridge with few nodes (small size hyperedge), as illustrated in Figure 2, algorithms that use hyperedge size tend to inappropriately contract bridge nodes (node 4 and node 7). This contraction can lead to the collapse of the overall structure of the hypergraph. Since the effective resistance of a bridge is high, algorithms based on effective resistance do not contract these nodes and preserve the integrity of the hypergraph structure.
Figure 2: The example of contraction in a simple hypergraph
@@ III-A Resistance-Based Hypergraph Clustering (Phase A)
III-A1 Hypergraph Clustering by Effective Resistances
A spectral hypergraph coarsening technique, HyperEF, effectively clusters nodes within each hyperedge that have low effective-resistance diameters as shown in Figure 3. This allows for the construction of significantly smaller hypergraphs while preserving the essential structural characteristics of the original hypergraph. A crucial element of HyperEF is an efficient method for estimating hyperedge effective resistances. This is accomplished by adapting the optimization-based effective-resistance estimation method shown in Eq. (6) to hypergraphs. Specifically, the effective resistance of a hyperedge is determined by finding an optimal vector $\chi^{*}$ through a designated optimization process.
@@ III-A1 Hypergraph Clustering by Effective Resistances
where $x^{\top}L_{G}x$ in Eq. (6) is replaced by the following nonlinear quadratic form $Q_{H}(\chi)$ [40]:
@@
HyperEF generates a significantly smaller hypergraph $H^{\prime}=(V^{\prime},E^{\prime},w^{\prime})$ from the original hypergraph $H=(V,E,w)$ by utilizing hyperedge effective resistances, achieving reductions in the numbers of vertices, edges, and weights ($|V^{\prime}|<|V|$, $|E^{\prime}|<|E|$, and $|w^{\prime}|<|w|$).
III-A2 Low-Resistance-Diameter Hypergraph Decomposition
Consider a weighted undirected graph $G=(\mathcal{V},\mathcal{E},z)$ where the weights $z\in\mathbb{R}^{\mathcal{E}}{>0}$ and $\gamma>1$ is sufficiently large. The effective-resistance diameter is defined as $\max\limits_{u,v\in\mathcal{V}}R_{eff}(u,v)$. Recent studies demonstrate that it is feasible to partition a simple graph $G$ into several node clusters $G[\mathcal{V}_{i}]$ that each have a low effective-resistance diameter, by removing only a constant fraction of the edges [53]:
@@ III-A2 Low-Resistance-Diameter Hypergraph Decomposition
Consider the weighted undirected graph $G=(\mathcal{V},\mathcal{E},z)$ where the weights $z\in\mathbb{R}^{\mathcal{E}}_{\geq 0}$, and let $\Phi_{G}$ represent the conductance of $G$. Cheeger’s inequality provides a way to establish a relationship between the effective-resistance diameter of the graph and its conductance [53]:
@@
The HyperEF algorithm is based on recent spectral hypergraph theory [40, 41] by extending the above theorem to the hypergraph setting: inequality (12) implies that it is possible to decompose a hypergraph into multiple (hyperedge) clusters that have small effective-resistance diameters by removing only a few inter-cluster hyperedges, while (13) implies that contracting the hyperedges (node clusters) with small effective-resistance diameters is not significantly impacting the original hypergraph conductance. HyperEF allows computing a much smaller hypergraph $H^{\prime}=(V^{\prime},E^{\prime},w^{\prime})$ given the original hypergraph $H=(V,E,w)$ by exploiting hyperedge effective resistances, which consists of the following three key steps:
Step 1
constructs the Krylov subspace to approximate the eigensubspace related to the original hypergraph;
Step 2
estimates the effective resistance of each hyperedge by applying the proposed optimization-based method;
Step 3
constructs the coarsened hypergraph by aggregating node clusters with low effective-resistance diameters.
@@ III-A3 Fast Estimation of Hyperedge Effective Resistances
To enhance efficiency, the search for $\chi^{*}$ in Eq. (10) is restricted to an eigensubspace represented by a few select Laplacian eigenvectors from the simplified graph derived from the original hypergraph. Consider $G_{b}=(\mathcal{V}_{b},\mathcal{E}_{b},z_{b})$ serves as the bipartite graph equivalent to the hypergraph $H=(V,E,w)$, where $|\mathcal{V}_{b}|=|V|+|E|$, $|\mathcal{E}_{b}|=\Sigma_{e\in E}|e|$, and $z_{b}$ denotes the scaled edge weights, specifically: $z(e,p)=\frac{w(e)}{d(e)}$. According to Eq. (6) computing approximate hyperedge effective resistances involves identifying a set of orthogonal eigenvectors that optimize (10).
@@
To avoid the computational complexity of determining eigenvalues and eigenvectors, we have used a scalable algorithm to approximate the eigenvectors by leveraging the Krylov subspace, defined as follows:
For a non-singular matrix $A_{n\times n}$, and a non-zero vector $\chi\neq 0\in\mathbb{R}^{n}$, the order-$(\rho+1)$ Krylov subspace generated by $A$ from $x$ is
@@
where $x$ is a random vector, and $A$ signifies the normalized adjacency matrix of the simple graph translated from the hypergraph via star expansion. Note that for each hypergraph, the Krylov subspace vectors in Eq. (14) only need to be computed once, which can be achieved in nearly linear time using only sparse matrix-vector operations. The effective resistance of each hyperedge can be estimated in constant time by identifying a few Krylov subspace vectors that maximize Eq. (10). As shown in Figure 4, the approximate effective resistances of simple graphs obtained using a few Krylov vectors can well correlate with the ground truths. Algorithm 1 provides the detailed flow of the proposed hyperedge effective resistance estimation method.
@@
It is assumed that $x^{(1)},x^{(2)},...,x^{(\rho)}\in\mathbb{R}^{\mathcal{V}_{b}}$ are the $\rho$ mutually-orthogonal vectors based on the order-$(\rho+1)$ Krylov subspace $\kappa_{\rho}(A,x)$, which is constructed in Step 1. The effective resistance estimation method, as described in (6), is extended by the HyperEF algorithm through the incorporation of the non-linear quadratic operator of hypergraphs (11), to include hypergraph spectral properties. By excluding the node embedding values associated with the star nodes from $x^{(i)}$, a new set of vectors $\chi^{(i)}$, all mutually orthogonal, is generated. Subsequently, each node in the hypergraph is embedded into a $\rho$-dimensional space. The resistance ratio ($r_{e}$) associated with a vector $\chi^{(i)}\in\chi^{(1)},...,\chi^{(\rho)}$ for each hyperedge $e$ is then computed as follows:
where $p$ and $q$ are identified as the two maximally-separated nodes in the $\rho$-dimensional embedding space. Multiple resistance ratios $r_{e}^{(1)},...,r_{e}^{(\rho)}$ are returned by Eq. (10), corresponding to $\chi^{(1)},...,\chi^{(\rho)}$.
After the resistance ratios are sorted in descending order, it is observed that:
$r_{e}^{1}>r_{e}^{2}>...>r_{e}^{\rho}.$ (16)
$R_{e}=\sum\limits_{i=1}^{m}r_{e}^{i},\quad e\in E.$ (17)
III-A4 Practical Implementations
To achieve good efficiency, the proposed spectral hypergraph coarsening framework leverages a linear-time local spectral embedding scheme based on low-pass filtering (smoothing) of random graph signals [54, 55, 25, 39, 56]. The key challenge of the above framework lies in designing universal low-pass filtering functions applicable to hypergraphs. To this end, low-pass graph filters for hypergraphs are designed based on the latest research in spectral hypergraph theory [57, 58, 50, 40]. One promising research direction is to convert hypergraphs into simple undirected graphs so that existing graph filters can be directly adopted for constructing the Krylov subspace required by the HyperEF algorithm. In addition, a spectrum-preserving edge expansion scheme is investigated based on: 1) clique expansion (CE), which replaces each hyperedge with multiple edges forming a clique among the incident vertices of the hyperedges; 2) star expansion (SE), which replaces each hyperedge with a new vertex connected to its incident vertices; 3) hybrid expansion (HE), which adaptively applies CE or SE. Specifically, the proposed spectral method determines whether a hyperedge should be replaced by a star or clique expansion according to the nonlinear diffusion operators defined on the hypergraph [40, 41], to achieve an optimal balance between complexity and quality. Compared with recent theoretical results in spectral methods for hypergraphs [50] that may not be efficiently implemented due to the polynomial-time algorithm complexity, the proposed spectral coarsening algorithms can be implemented using parallel-friendly sparse kernels, which is suitable for accelerations on modern hardware platforms, such as CPUs, GPUs, and FPGAs [59, 60, 61, 62, 63, 64, 65, 66].
III-A5 Multilevel Effective Resistance Clustering
@@
We enhance the hypergraph structure by employing a technique that iteratively computes a vector of effective resistance $R$ and contracts hyperedges exhibiting low effective resistances ($R_{e}<\delta$), where $\delta$ is the effective resistance threshold. Throughout this process, node clusters are contracted by merging nodes within each cluster and replacing them with a new entity termed a supernode at the subsequent level. The structural information of the hypergraph is preserved and transmitted through levels by assigning weights to the supernodes, which are equal to the hyperedge effective resistances evaluated at the preceding level.
Let $H^{(l)}=(V^{(l)},E^{(l)},w^{(l)})$ represent the hypergraph at the $l$-th level, where $S\in V^{(l)}$ is a cluster of nodes producing a supernode $\vartheta\in V^{(l+1)}$ at level $l+1$. The vector of node weights is denoted as $\eta^{(l)}\in\mathbb{R}_{\geq 0}^{V^{(l)}}$, initially set to all zeros for the original hypergraph. At each level $l$, $\eta$ is updated by:
$\eta_{\vartheta}:=\sum\limits_{j=1}^{|S|}\eta(v_{j}^{(l)}).\quad v\in V^{(l)}:v\in S$ (18)
III-A6 Node Weight Propagation (NWP)
The vector of effective resistance $R$ is updated based on the node weights $\eta$ to transmit the clustering information from previous levels to the current level:
$R_{e}^{(l)}=\sum\limits_{k=1}^{|e|}\eta(v_{k}^{(l)})+R_{e}^{(l)}.$ (19)
Consequently, the effective resistance of a hyperedge at a coarser level depends not only on the effective resistance ($R_{e}^{(l)}$) evaluated at the current level, but also on the data transferred from all previous (finer) levels.
@@ III-A6 Node Weight Propagation (NWP)
The experimental results indicate that using Eq. (19) for effective resistances estimation yields more balanced hypergraph clustering outcomes compared to approaches that ignores the previous clustering information. The complete workflow for the effective resistance clustering algorithm, HyperEF, used for coarsening a hypergraph $H$ across $L$ levels, is detailed in Algorithm 2.
@@
4:    Compute the node weights using (18).
5:    Update the effective resistance vector $R$ by applying (19).
6:    Sort the hyperedges with ascending $R$ values.
7:    Starting with the hyperedges that have the lowest effective resistances, contract (cluster) the hyperedge (nodes) if $R_{e}<\delta$.
8:    Construct a coarsened hypergraph $H^{\prime}$ accordingly.
9:  end for
10:  Return $H^{\prime}$.

### v3

@@ II-A5 Effective resistance distance
Graph conductance is a metric used to evaluate the quality of connectivity within a subset of nodes in a graph. Specifically, a lower conductance value indicates that the subset exhibits a higher number of internal edges relative to the number of edges connecting it to the remainder of the graph. This suggests that the subset forms a strongly-connected cluster with limited external interaction. Alternatively, when a graph is modeled as a resistive electrical network, where nodes represent junctions and edge weights correspond to resistances, the concept of effective resistance between two nodes serves as a measure of their pairwise connectivity. A smaller effective resistance implies the presence of multiple alternative paths between the nodes, indicating stronger connectivity in the network.
@@ II-A7 Hypergraph conductance
A hypergraph $H=(V,E,w)$ consists of a vertex set $V$ and a set of hyperedges $E$ with weights $w$. The degree of a vertex $d_{v}$ is defined as: $d_{v}:=\Sigma_{e\in E:v\in e}w(e)$, where $w(e)$ represents the weight of each hyperedge. The volume of a node set $S\subseteq V$ in the hypergraph is defined as: $vol(S):=\Sigma_{v\in S}d_{v}$. The conductance of a subset $S$ within the hypergraph is then calculated as:
@@ II-B3 Partitioning objectives
Hypergraph partitioning extends the concept of graph partitioning. Its objective is to distribute the vertex set into multiple disjoint subsets while minimizing a specified cut metric and adhering to certain imbalance constraints. The process of dividing into two subsets is known as bipartitioning, whereas dividing into multiple subsets, typically referred to as $k$-way partitioning, involves partitioning into $k$ parts. More formally, consider a hypergraph $H=(V,E,w)$, where $k$ is a positive integer (with $k\geq 2$) and $\epsilon$ is a positive real number (where $\epsilon\leq\frac{1}{k}$). The objective of $k$-way balanced hypergraph partitioning is to divide $V$ into $k$ disjoint subsets $S=\{V_{0},V_{1},\dots,V_{k-1}\}$ such that:
@@ III SHyPar: Hypergraph Partitioning via Spectral Coarsening
To address the limitations of existing hypergraph coarsening methods that rely on simple heuristics, we propose a theoretically sound and practically efficient framework for hypergraph coarsening. Specifically, we propose a two-phase spectral hypergraph coarsening scheme based on the recent research on spectral hypergraph clustering [46, 47]. Phase A utilizes spectral hypergraph coarsening (HyperEF) to decompose a given hypergraph into smaller node partitions with bounded effective-resistance diameters [47]. This is followed by Phase B, which guides the coarsening stage using a flow-based community detection method (HyperSF) aimed at minimizing the ratio cut [46]. Next, we exploit the proposed two-phase spectral hypergraph coarsening method for multilevel hypergraph partitioning: the prior heuristic hypergraph coarsening schemes will be replaced by the proposed spectral coarsening methods to create a hierarchy of coarser hypergraphs that can preserve the key structural properties of the original hypergraph.
III-A Resistance-Based Hypergraph Coarsening (Phase A)
We leverage the effective resistance parameter to coarsen the hyperedges successively by contracting the hyperedges with small effective resistance. Effective resistance in simple graphs has been used to detect the critical edges for the global structure and it shows how different graph components are connected with each other.
Limitations of Existing Coarsening Methods
The existing coarsening algorithms contract vertices at each level of the hierarchy. The primary method involves contracting each vertex with the best neighboring node. This is commonly done by using rating functions to identify and contract highly connected vertices. However, these determinations often rely solely on the weights and sizes of the hyperedges. Rating functions, such as those described in Eq. (9), based on hyperedge size are limited to the local structural properties of the hypergraph and do not account for its global structure. In contrast, the effective-resistance diameter provides a more comprehensive criterion, which considers the global structure. To illustrate this limitation, consider a scenario where the hypergraph contains a bridge with few nodes (small size hyperedge), as illustrated in Figure 2. Algorithms that use hyperedge size tend to inappropriately contract bridge nodes (node 4 and node 7). This contraction can lead to the collapse of the overall structure of the hypergraph. Since the effective resistance of a bridge is high, algorithms based on effective resistance do not contract these nodes and preserve the integrity of the hypergraph structure.
Figure 2: The example of contraction in a simple hypergraph.
@@ III-A Resistance-Based Hypergraph Coarsening (Phase A)
III-A1 HyperEF: Hypergraph Coarsening via effective resistance Clustering
A spectral hypergraph coarsening technique, HyperEF, clusters nodes within each hyperedge when those nodes have a low effective-resistance diameter, as illustrated in Figure 3. This approach significantly reduces the hypergraph size while preserving the original structural characteristics. A crucial part of HyperEF is an efficient procedure for estimating hyperedge effective resistances, which adapts the optimization-based method from Eq. (6) to hypergraphs. Specifically, the effective resistance of a hyperedge is determined by finding an optimal vector $\chi^{*}$ via the following optimization:
@@ III-A1 HyperEF: Hypergraph Coarsening via effective resistance Clustering
where the original quadratic form $x^{\top}L_{G}x$ in Eq. (6) is replaced by the nonlinear quadratic form $Q_{H}(\chi)$ [40]:
@@
As shown in Figure 3, HyperEF generates a significantly smaller hypergraph $H^{\prime}=(V^{\prime},E^{\prime},w^{\prime})$ from the original hypergraph $H=(V,E,w)$ utilizing effective resistances of the hyperedge, achieving reductions in the number of vertices, edges and weights ($|V^{\prime}|<|V|$, $|E^{\prime}|<|E|$ and $|w^{\prime}|<|w|$).
III-A2 Low-Resistance-Diameter Decomposition
Consider a weighted undirected graph $G=(\mathcal{V},\mathcal{E},z)$ with weights $z\in\mathbb{R}^{\mathcal{E}}{>0}$ and sufficiently large $\gamma>1$. The effective-resistance diameter is defined as $\max\limits_{u,v\in\mathcal{V}}R_{eff}(u,v)$. Recent studies demonstrate that it is possible to partition a graph $G$ into multiple node clusters $G[\mathcal{V}_{i}]$ with low effective-resistance diameters by removing only a small fraction of edges [53]:
@@ III-A2 Low-Resistance-Diameter Decomposition
Additionally, let $\Phi_{G}$ represent the conductance of $G$. Cheeger’s inequality provides a way to establish a relationship between the effective-resistance diameter of the graph and its conductance [53]:
@@
Leveraging recent spectral hypergraph theory [40, 41], HyperEF extends the above theorems to hypergraphs. Inequality (12) implies that it is possible to decompose a hypergraph into multiple (hyperedge) clusters that have small effective-resistance diameters by removing only a few inter-cluster hyperedges, while (13) implies that contracting the hyperedges (node clusters) with small effective-resistance diameters will not significantly impact the original hypergraph conductance. Based on these theoretical foundations, HyperEF consists of the following three main steps:
  • Constructing the Krylov subspace to approximate the eigensubspace related to the original hypergraph;
  • Estimating the effective resistance of each hyperedge by applying the proposed optimization-based method;
  • Constructing the coarsened hypergraph by aggregating node clusters with low effective-resistance diameters.
@@ III-A3 Fast Estimation of Hyperedge Effective Resistances
To achieve high efficiency, the search for $\chi^{*}$ in Eq. (10) is restricted to an eigensubspace represented by a few select Laplacian eigenvectors from the simplified graph derived from the original hypergraph. Consider $G_{b}=(\mathcal{V}_{b},\mathcal{E}_{b},z_{b})$ as the bipartite graph equivalent to the hypergraph $H=(V,E,w)$, where $|\mathcal{V}_{b}|=|V|+|E|$, $|\mathcal{E}_{b}|=\Sigma_{e\in E}|e|$, and $z_{b}$ denotes the scaled edge weights: $z(e,p)=\frac{w(e)}{d(e)}$.
According to Eq.(6), hyperedge effective resistances can be approximated by identifying a set of orthogonal eigenvectors that maximize Eq. (10).
@@
To avoid the computational complexity of determining eigenvalues and eigenvectors, a scalable algorithm is used in HyperEF to approximate the eigenvectors by leveraging the Krylov subspace, defined as follows:
For a non-singular matrix $A_{n\times n}$ and a non-zero vector $\chi\neq 0\in\mathbb{R}^{n}$, the order-$(\rho+1)$ Krylov subspace generated by $A$ from $x$ is:
@@
where $x$ is a random vector and $A$ is the normalized adjacency matrix of the simple graph obtained from the hypergraph via star expansion.
Let $x^{(1)},x^{(2)},...,x^{(\rho)}\in\mathbb{R}^{\mathcal{V}_{b}}$ be the $\rho$ mutually-orthogonal vectors based on the order-$(\rho+1)$ Krylov subspace $\kappa_{\rho}(A,x)$. The effective resistance estimation in Eq. (6) is extended in HyperEF by incorporating the non-linear quadratic operator of hypergraphs from Eq. (11) to include the spectral properties of the hypergraph. By excluding the node embedding values associated with the star nodes in $x^{(i)}$, we generate a new set of vectors $\chi^{(i)}$ that are all mutually orthogonal. Each node in the hypergraph can then be embedded into a $\rho$-dimensional space. The resistance ratio ($r_{e}$) associated with a vector $\chi^{(i)}\in\chi^{(1)},...,\chi^{(\rho)}$ for each hyperedge $e$ is computed as:
where $p$ and $q$ are the two maximally-separated nodes in the $\rho$-dimensional embedding space. Multiple resistance ratios $r_{e}^{(1)},...,r_{e}^{(\rho)}$ are returned by Eq. (10), corresponding to $\chi^{(1)},...,\chi^{(\rho)}$. After sorting resistance ratios in descending order, we have:
$r_{e}^{1}>r_{e}^{2}>...>r_{e}^{\rho}.$ (16)
$R_{e}=\sum\limits_{i=1}^{m}r_{e}^{i},\quad e\in E.$ (17)
Note that for each hypergraph, the Krylov subspace vectors in Eq. (14) only need to be computed once, which can be achieved in nearly linear time using only sparse matrix-vector operations. The effective resistance of each hyperedge can be estimated in constant time by identifying a few ($m$) Krylov subspace vectors that maximize the resistance ratio in Eq. (15). As shown in Figure 4, the approximate effective resistances of simple graphs obtained using a few Krylov vectors correlate well with the ground-truth values, with a correlation coefficient of about 0.76. Algorithm 1 provides the detailed flow of the proposed hyperedge effective resistance estimation method.
@@
III-A4 Hierarchical Effective Resistance Propagation
@@
We preserve the hypergraph structure by employing a technique that iteratively computes a vector of effective resistance $R$ and contracts hyperedges exhibiting low effective resistances ($R_{e}<\delta$), where $\delta$ is the effective resistance threshold. Throughout this process, node clusters are contracted by merging the nodes within each cluster and replacing that cluster with a new supernode in the next level. By assigning a weight to each supernode—equal to the hyperedge’s effective resistance evaluated at the previous level—we propagate essential structural information through all levels of coarsening. As an example, in Figure 3, when hyperedge $e_{4}$ in $H$ is contracted, the node 8 in $H^{\prime}$ takes on a weight that is equal to the effective resistance of $e_{4}$. This ensures that the structural information from the original hyperedge remains available in subsequent coarsening steps.
Let $H^{(l)}=(V^{(l)},E^{(l)},w^{(l)})$ represent the hypergraph at the $l$-th level. The vector of effective resistance $R$ is updated in each coarsening level according to the following equation:
$R_{e}^{(l)}\leftarrow\sum_{v\in e}\eta(v)+R_{e}^{(l)},$ (18)
where $\eta(v)$ represents the weight of the nodes $v\in e$ corresponding to a contracted hyperedge from the previous level, initially set to all zeros for the original hypergraph. Consequently, the effective resistance of a hyperedge at a coarser level depends not only on $R_{e}^{(l)}$ (computed at the current level), but also on the effective-resistance data transferred from all previous levels. This allows the algorithm to preserve global structural information through the multilevel coarsening process.
@@
The experimental results indicate that using Eq. (18) for effective resistances estimation yields more balanced hypergraph clustering outcomes compared to approaches that ignore the previous clustering information. The complete workflow for the effective resistance clustering algorithm, HyperEF, used for coarsening a hypergraph $H$ across $L$ levels, is detailed in Algorithm 2.
@@
4:    Update the effective resistance vector $R$ by including the supernode weights using Eq. (18)
5:    Sort the hyperedges with ascending $R$ values.
6:    Starting with the hyperedges that have the lowest effective resistances, contract (cluster) the hyperedge (nodes) if $R_{e}<\delta$.
7:    Construct a coarsened hypergraph $H^{\prime}$ accordingly.
8:  end for
9:  Return $H^{\prime}$.